# Evaluating diversion and treatment policies for opioid use disorder

### v3

@@ A.5.6 Distribution (6) - Time in the active state given next event is stop opioid use
The quantile $q$ was estimated from Rivera et al., (2018), where they estimated that the mean confidence interval for the proportion of individuals developing heroin dependence was between .167 and .506 in 2016. This year was chosen because it aligned with the time frame from most of the dataset. We assume that if someone fails to develop dependence, this could alternatively be interpreted as the individual stopping opioid use. Therefore, we can use the range $(1-0.506,1-0.167)$ to estimate the proportion of individuals that move to the inactive use state. To estimate the $z_{q}$, or the timeframe individuals move to inactive use, we used Bauer, (2019), who concluded that “heroin problems and experiences now begin to coalesce into heroin use disorder syndrome within 90-120 days after first heroin use”. During model calibration, we tested a baseline, low, and high mix of these ranges. We ended up with our high estimate where we estimated about 49.4% of individuals would be in the inactive use state within 120 days after using an opioid since moving to the active use state. This resulted in a $\gamma=0$, $m=1$, and $z_{0.494}=120$ days to obtain a $\mu_{6}=4.82$ and $\sigma_{6}=2.20$.
@@ A.5.7 Distribution (7) - Time in the active state given next event is non-opioid related death
Distribution (7) is an empirical distribution of the length of remaining life based on the individual’s current age. Age estimation is described in Section A.4. We then sample the time to death from an empirical distribution estimated from “Table 1. Abridged life tables for all causes of death combined and eliminating specified causes, for the total population: United States, 1999-2001” from Arias et al., (2013). The distribution is rescaled depending on the current individual’s age to obtain a final time until non-opioid-related death.
@@ A.6.1 Distribution (A) - Time in the hosptital or ED
Distribution (A) parameters were estimated using the study done by Singh and Cleveland, (2020), which analyzed U.S. National Inpatient Sample (NIS) data from 1998–2016 to estimate OUD hospitalization rates. OUD-hospitalization characteristics by the U.S. hospital region report a median length of stay in the Midwest (total Midwest hospital stays was 182,787), where Dane County is located, of 1.8 days. Additionally, they report 72.3% of patients were in the hospital for less than or equal to $3$ days. Therefore, we estimated our model parameters using $\gamma=0$, $m=1.8$, and $z_{0.723}=3$ to obtain $\mu_{A}=0.82$ and $\sigma_{A}=0.48$.
@@ A.6.2 Distribution (B) - Time in the CJS
Distribution (B) parameters were estimated directly from the dataset collected apart of the study done by Zhang et al., (2022), which compares police contacts, arrests, and incarceration between a historical control group ($N=52$) and an arrest diversion group ($N=263$). The data showed that for the historical comparison group, 47 of the individuals spent less than 57 days in jail after being arrested for an offense that would have qualified them for an illicit drug arrest diversion program. The mode was one day. Therefore, we estimated our model parameters using $\gamma=0$, $m=1$, $z_{0.9}=57$ days to obtain $\mu_{B}=2.16$ and $\sigma_{B}=1.47$. The same estimates were used time in the CJS due to non-drug-related crime for two reasons. First, among the 79 non-opioid related crimes 0 to 12 months before the arrest diversion eligible crime, 75 of the arrests resulted in spending less than 91 days in jail, with a mode of one day, which is well within the range considered for distribution (B) in our sensitivity analysis. Secondly, limiting the number of estimated distributions reduces model complexity.
@@ A.6.3 Distribution (C) - Time in OUD treatment
Distribution (C) parameters were estimated from Table 51 in Division of Care and Teatment Services, (2018), which indicates the percentage of patients receiving at least 90 days of any SUD treatment in Dane County was 59.5%. The most likely value, $m$, is selected through model calibration by adding approximately ±$\pm 5$% from a starting value of 30. Therefore, we estimated our model parameters using $\gamma=0$, $m=30$, and $z_{1-.595}=90$ days to obtain $\mu_{C}=4.78$ and $\sigma_{C}=1.18$.
@@ A.6.4 Distribution (D) - Time in the inactive state given the previous state was the CJS
Distribution (D) parameters were estimated from two separate studies. In Table 2 of Kinlock et al., (2008), $\nicefrac{{48}}{{62}}=77.4$% of males who were released from prison or jail in Baltimore and received counseling only for their opioid use during prison or jail had used heroin within 90 days from release. In Figure 1 from Bukten et al., (2017), individuals ($N=91,090$) from the 15-year cohort study who were released from prison in Norway were most likely to relapse (and die of an overdose death) on the second day. Therefore, we estimated our model parameters using $\gamma=0$, $m=2$, and $z_{0.774}=90$ days to obtain $\mu_{D}=3.29$ and $\sigma_{D}=1.61$. These estimates were used for opioid and non-opioid-related arrests that led to time in the CJS.
@@ A.6.5 Distribution (E) - Time in the inactive state given the previous state was OUD treatment
Distribution (E) parameters were estimated from Nunes et al., (2018), where the study had two types of treatment, naltrexone, and treatment of OUD as usual. It was further split into long and short inpatient and outpatient care. The most likely value, $m$, was estimated from Figure 1 using the outpatient and short-term treatment as usual groups, where the fourth week (i.e., 28 days) had the largest drop in treatment retention. We also estimated from Table 3 the cumulative opioid use at the end of 26 weeks (i.e., 182 days) that $\nicefrac{{226}}{{308}}=0.734$ of patients in all treatment groups had used an opioid within 182 days. Therefore, we estimated our model parameters using $\gamma=0$, $m=28$, $q=0.734$, and $z_{0.734}=182$ days to obtain $\mu_{E}=4.52$ and $\sigma_{E}=1.09$.
@@ A.6.6 Distribution (F) - Time in the inactive state given the previous state was the hospital or ED
Distribution (F) parameters were estimated from Chutuape et al., (2001), where 21.3% of individuals ($N=116$) self-reported not using after 30 days following release of a three-day inpatient opioid detoxification. However, the researchers reported that urine samples indicate this is closer to 12-15%, as stated on page 31. The most likely value, $m$, is informed from Figure 2, where 26% of 66 individuals reported using heroin on the day of discharge. Therefore, we estimated our model parameters using $\gamma=0$, $m=1$, and $z_{0.85}=30$ days to obtain $\mu_{F}=1.95$ and $\sigma_{F}=1.40$.
@@ A.6.7 Distribution (G) - Time in the inactive state given the previous state was active opioid use
Distribution (G) parameters were calibrated. The quantile $z_{q}$ was selected through model calibration by adding approximately $\pm$5from a starting value assessment that 90% of individuals used opioids within 20 years. The most likely value $m$ is selected through model calibration by adding approximately $\pm$5from a starting value 30. The final model parameters were estimated using $\gamma=0$, $m=1$, and $z_{0.7}=5.5*365.25$ days per year to obtain $\mu_{G}=6.29$ and $\sigma_{G}=2.51$.
@@ Appendix B Determining Number of Replications
$\hat{s}$ at $n=600$ 262.34 10.58 28.48 57.18 60.41
$h$ at $n=600$ 21.99 0.86 2.21 4.82 4.55
$\hat{n}$ 27 432 126 505 563
Additionally, we checked the residuals of the main outputs for normality by plotting and observing Q-Q plots and performing K-S tests of the main outputs. Table 8 reports the K-S test with Lilliefors correction and skewness. The K-S test fails to reject the null hypothesis that the number of arrests, number of hospital encounters, and number of treatments in 2032 follow a normal distribution. The K-S test rejects the null hypothesis for the number of individuals in the active use state, the inactive use state, and the number of opioid-related deaths at the end of 2032. Looking at the skewness in Table 8 we see that the active and inactive use state skews slightly positively, meaning the mean estimates used in the main analysis are above the median of all $n$ runs. Similarly, the number of opioid-related deaths is skewed negatively, meaning the mean used in the analysis is lower than the median of the number of $n$ runs. Figure 4 shows the Q-Q plots for the five main outputs of the model for years 2016, 2018, and 2032. Based on the relatively small skewness ($<0.75$) and Q-Q plots we conclude that all the main outputs are sufficiently normal to use confidence intervals as the main results, yet keep skewness in mind when interpreting the results. We also note that the normality assumption is valid for our cumulative results since the sum of two jointly normal random variables also follows a normal distribution.In fact, following the same process as above with cumulative outputs, showed the cumulative outputs required fewer than 400 replications to achieve a 95% CI. We have omitted the cumulative results due to their similarity to the yearly outputs.
@@
K-S Testα (p-value) $<$0.001 0.061 0.159 0.374 0.404
Skewness 0.305 -0.110 -0.005 0.175 0.158
@@ Appendix C Additional Simulated Results
Figure 7 illustrates the mean 95% CIs of arrests, hospitalizations, and treatment episodes per person over time. One is the lower bound of the re-arrest, re-hospitalization, and OUD treatment re-start rates. At one, everyone who used the given system only used it once in a given year. We expect the base model re-arrest and re-hospitalization rates to be a lower bound on the number of interactions for a given individual with these systems. In Figure 7, we see statistically significant decreases in re-hospitalizations for policies with increases in OD, decreases in re-arrests for policies with AD or CM, and increases in OUD treatment for all policies in the first year or two after policy implementation. Figure 7 shows similar results to Table 5, but over time.
@@ Appendix D Sensitivity Analysis
This section explores and describes the sensitivity of the DES model. We conducted the sensitivity analysis on the scenario with arrest diversion at 60%, overdose diversion at 80%, and case management at 60%, referred to as scenario (60, 80, 60) for the remainder of this section. This scenario was selected for two main reasons. The first was due to the scenario’s high percentages of individuals diverted by the interventions. Two, the scenario’s scale of implementation is feasible for stakeholders to implement since we can expect some eligible individuals not to complete or participate in an available program (White et al.,, 2021). We evaluate changes to the following outputs in 2016, 2018, and 2032: number of overdose deaths, individuals entering the CJS, hospital encounters, treatment, and active use. These outputs were selected since they were the main focus of the results. The outputs in the year 2016 test input sensitivity to the base model when no interventions have been implemented yet. The year 2018 indicates the sensitivity of the model with increased overdose diversion and arrest diversion implemented, and the year 2032 indicates the effects from the results still hold and the model sensitivity to when all three interventions are implemented.
@@
1 Warmup Initiation Age, $\mu$ 2.88 (2.73, 3.02)
2 Warmup Initiation Age, $\sigma$ 0.56 (0.53, 0.59)
3 Warmup Prevalence Age, $\mu$ 3.10 (2.95, 3.26)
4 Warmup Prevalence Age, $\sigma$ 0.46 (0.44, 0.48)
@@
Table 11 indicates three very strong negative monotonic relationships between the inputs and outputs. The first is the relationship between the number of arrests and parameter 17, $\mu$ of the distribution for the time in the active state given the next event is an opioid-related arrest. The second is the relationship between the number of hospital encounters and parameter 15, $\mu$ of the distribution for the time in the active state given that the next event is an opioid-related hospital encounter in 2018 and 2032, with strong monotonicity in 2016. The third is the relationship between the number of OUD treatment starts and parameter 19, $\mu$ of the distribution for the time in the active state given the next event is OUD treatment in the year 2016, with strong monotonicity in 2018 and moderate monotonicity in the year 2032. This relationship between OUD treatment and parameter 19 is less monotonous after implementing the OUD policies. As shown in Table 12, all three very strong monotonicity relationships have effect sizes of -8.458, -5.035, and -6.326, respectively, the greatest of which is that of the relationship of parameter 17 and the number of opioid-related arrests. It is also worth noting the corresponding $\sigma$ parameters of these LN distributions had notable OLS effects of 3.12, 1.669, and 0.578, respectively, even though the relationships were weakly monotone or determined non-monotone. This suggests that the corresponding model outputs are more sensitive to small changes to the $\mu$ parameters over the $\sigma$ parameters. Additionally, the adjustments are more predictable (i.e., linear) to the corresponding output with changes to the $\mu$ parameters. Therefore, parameters 15 through 20 should be carefully selected as they directly impact the magnitude of outputs over time. We do note that the effect size of parameter 15 on hospital encounters and parameter 18 on OUD treatment starts to decrease after the sustained implementation of the policies. This means that while they still impact model outputs, they seem to affect the outputs less after the policies are implemented.
@@
The other strong monotone relationships in the model are between the number of opioid-related deaths and parameters 11 and 15. Parameter 11 is the $\mu$ of the distribution for the active state time given that the next event is an opioid-related death pre-2019. The corresponding $\sigma$ parameter, 12, is also moderately monotone for this distribution. We note that the relationship between opioid-related deaths and parameter 14, the $\mu$ of the distribution for the active state time given that the next event is an opioid-related death pre-2019, is weakly monotone with a smaller effect size than that of parameter 11. Since the model is sensitive to parameters 11 and 12, care should be given to parameterizing opioid-related deaths. We also note that after the implementation of AD and OD, the absolute value of the OLS effects of parameters 11 and 12 reduces the magnitude from -5.752 to -1.064 and 3.625 to 0.646, respectively. This suggests that the AD and OD interventions make the model less sensitive to parameter 11. Parameter 15 is the $\mu$ of the distribution for the active state time given that the next event is a hospital encounter. Parameter 15 becomes more monotone, and its effect size increases after the AD and OD policies are implemented in 2018 but is decreased for 2032. This suggests that as parameter 15 increases, an OD policy would look more competitive for opioid-related deaths. The $\mu$ parameters 11, 13, and 15 have larger effect sizes than their corresponding distribution’s $\sigma$ parameters, meaning the model is more sensitive to the $\mu$ parameters of the distribution than their $\sigma$ parameters.
@@
Several additional moderate to weak monotone relationships exist among parameters 5, 10, 21, and 40. As the size of the total starting population (i.e., parameter 5) increases, the number of individuals in the active use state at the end of 2032 moderately increases. Parameter 5 also has a weakly positive monotone relationship with overdose-related deaths in 2016 and 2018 with effect sizes of 0.415 and 0.297, opioid-related arrests in 2016 with an effect size of 0.375, opioid-hospital encounters in 2016 and 2018 with effect sizes of 0.322 and 0.302, and OUD treatment in 2016 and 2018 with effect sizes of 0.387 and 0.316. This suggests that the starting population slightly affects all states, with the greatest impact on the active use state, but that its influence reduces the longer the simulation is run. As the opioid initiation arrival rate weakly increases (i.e., parameter 10), so does the number of individuals in the active use state with an effect size of 0.861 in 2032.
Parameter 24, $\mu$ of the distribution for the time in the active state given the next event is to sop opioid use, is moderately monotone, with the number of active use with an effect size of 2.5385 in 2016, and OUD treatments with an effect size of 2.326 in 2018. Parameter 24 also has a weak positive monotone relationship with the number of opioid-related deaths, opioid-related arrests, and hospital encounters, with effect sizes of 1.817, 2.28, and 1.719, respectively. Lastly, the $\mu$ of the LN distribution for the time in the inactive state following the active state (i.e., parameter 40) has a moderate monotone relationship with the number of active use starts and the number of OUD treatments in the year 2032, with effect sizes of -2.446 and -2.028, respectively. Parameter 38 is weakly monotone with d OUD treatment starts in 2016 and 2018, opioid-related deaths, opioid-related arrests, and opioid-related hospital encounters. This suggests that the model is highly sensitive to parameter 38.
@@
The monotonic relationships between the input and output parameters described in this section are expected based on the model structure. Additionally, except for opioid-related deaths, the effect size of parameters on outputs does not change as policies are implemented. This suggests that changes to these parameters would not affect the conclusions of this paper between the base model and treatment policy scenarios regarding opioid-related arrests, hospital encounters, treatments, and active use.

### v4

@@ A.5.6 Distribution (6) - Time in the active state given next event is stop opioid use
The quantile $q$ was estimated from Rivera et al., (2018), where they estimated that the mean confidence interval for the proportion of individuals developing heroin dependence was between .167 and .506 in 2016. This year was chosen because it aligned with the time frame from most of the dataset. We assume that if someone fails to develop dependence, this could alternatively be interpreted as the individual stopping opioid use. Therefore, we can use the range $(1-0.506,1-0.167)$ to estimate the proportion of individuals that move to the inactive use state. To estimate the ${x}_{q}$, or the timeframe individuals move to inactive use, we used Bauer, (2019), who concluded that “heroin problems and experiences now begin to coalesce into heroin use disorder syndrome within 90-120 days after first heroin use”. During model calibration, we tested a baseline, low, and high mix of these ranges. We ended up with our high estimate where we estimated about 49.4% of individuals would be in the inactive use state within 120 days after using an opioid since moving to the active use state. This resulted in a $\gamma=0$, $m=1$, and ${x}_{0.494}=120$ days to obtain a $\mu_{6}=4.82$ and $\sigma_{6}=2.20$.
@@ A.5.7 Distribution (7) - Time in the active state given next event is non-opioid related death
Distribution (7) is an empirical distribution of the length of remaining life based on the individual’s current age. Age estimation is described in Section A.4. We use “Table 1. Abridged life tables for all causes of death combined and eliminating specified causes, for the total population: United States, 1999-2001” from Arias et al., (2013) to estimate the number of deaths that occur in each age interval when eliminating deaths due to Drug-induced causes (i.e., death codes F11.0–F11.5, F11.7–F11.9, F12.0–F12.5, F12.7–F12.9, F13.0–F13.5, F13.7–F13.9, F14.0–F14.5, F14.7–F14.9, F15.0–F15.5, F15.7–F15.9, F16.0–F16.5, F16.7–F16.9, F17.0, F17.3–F17.5, F17.7–F17.9, F18.0–F18.5, F18.7–F18.9, F19.0–F19.5, F19.7–F19.9, X40–X44, X60–X64, X85, and Y10–Y14). From there, we create a survival function using the Kaplan-Meier Estimator (Kaplan and Meier,, 1958), implemented through the Python KaplanMeierFitter function as part of the lifelines package (Davidson-Pilon,, 2024). Therefore, we estimate the expected age at which an individual will experience a non-opioid-related death and the number of days until that death from when the individual enters the simulation, given the current simulation time and the estimated current age of the individual.
@@ A.6.1 Distribution (A) - Time in the hosptital or ED
Distribution (A) parameters were estimated using the study done by Singh and Cleveland, (2020), which analyzed U.S. National Inpatient Sample (NIS) data from 1998–2016 to estimate OUD hospitalization rates. OUD-hospitalization characteristics by the U.S. hospital region report a median length of stay in the Midwest (total Midwest hospital stays was 182,787), where Dane County is located, of 1.8 days. Additionally, they report 72.3% of patients were in the hospital for less than or equal to $3$ days. Therefore, we estimated our model parameters using $\gamma=0$, $m=1.8$, and ${x}_{0.723}=3$ to obtain $\mu_{A}=0.82$ and $\sigma_{A}=0.48$.
@@ A.6.2 Distribution (B) - Time in the CJS
Distribution (B) parameters were estimated directly from the dataset collected apart of the study done by Zhang et al., (2022), which compares police contacts, arrests, and incarceration between a historical control group ($N=52$) and an arrest diversion group ($N=263$). The data showed that for the historical comparison group, 47 of the individuals spent less than 57 days in jail after being arrested for an offense that would have qualified them for an illicit drug arrest diversion program. The mode was one day. Therefore, we estimated our model parameters using $\gamma=0$, $m=1$, ${x}_{0.9}=57$ days to obtain $\mu_{B}=2.16$ and $\sigma_{B}=1.47$. The same estimates were used time in the CJS due to non-drug-related crime for two reasons. First, among the 79 non-opioid related crimes 0 to 12 months before the arrest diversion eligible crime, 75 of the arrests resulted in spending less than 91 days in jail, with a mode of one day, which is well within the range considered for distribution (B) in our sensitivity analysis. Secondly, limiting the number of estimated distributions reduces model complexity.
@@ A.6.3 Distribution (C) - Time in OUD treatment
Distribution (C) parameters were estimated from Table 51 in Division of Care and Teatment Services, (2018), which indicates the percentage of patients receiving at least 90 days of any SUD treatment in Dane County was 59.5%. The most likely value, $m$, is selected through model calibration by adding approximately ±$\pm 5$% from a starting value of 30. Therefore, we estimated our model parameters using $\gamma=0$, $m=30$, and ${x}_{1-.595}=90$ days to obtain $\mu_{C}=4.78$ and $\sigma_{C}=1.18$.
@@ A.6.4 Distribution (D) - Time in the inactive state given the previous state was the CJS
Distribution (D) parameters were estimated from two separate studies. In Table 2 of Kinlock et al., (2008), $\nicefrac{{48}}{{62}}=77.4$% of males who were released from prison or jail in Baltimore and received counseling only for their opioid use during prison or jail had used heroin within 90 days from release. In Figure 1 from Bukten et al., (2017), individuals ($N=91,090$) from the 15-year cohort study who were released from prison in Norway were most likely to relapse (and die of an overdose death) on the second day. Therefore, we estimated our model parameters using $\gamma=0$, $m=2$, and ${x}_{0.774}=90$ days to obtain $\mu_{D}=3.29$ and $\sigma_{D}=1.61$. These estimates were used for opioid and non-opioid-related arrests that led to time in the CJS.
@@ A.6.5 Distribution (E) - Time in the inactive state given the previous state was OUD treatment
Distribution (E) parameters were estimated from Nunes et al., (2018), where the study had two types of treatment, naltrexone, and treatment of OUD as usual. It was further split into long and short inpatient and outpatient care. The most likely value, $m$, was estimated from Figure 1, where the fourth week (i.e., 28 days) had the largest drop in treatment retention when considering all treatment groups. This is driven by the outpatient and short-term treatment groups, noting that the outpatient group is approximately four times the size of the short-term and long-term treatment groups. We also estimated from Table 3 the cumulative opioid use at the end of 26 weeks (i.e., 182 days) that $\nicefrac{{226}}{{308}}=0.734$ of patients in all treatment groups had used an opioid within 182 days. Therefore, we estimated our model parameters using $\gamma=0$, $m=28$, $q=0.734$, and ${x}_{0.734}=182$ days to obtain $\mu_{E}=4.52$ and $\sigma_{E}=1.09$.
@@ A.6.6 Distribution (F) - Time in the inactive state given the previous state was the hospital or ED
Distribution (F) parameters were estimated from Chutuape et al., (2001), where 21.3% of individuals ($N=116$) self-reported not using after 30 days following release of a three-day inpatient opioid detoxification. However, the researchers reported that urine samples indicate this is closer to 12-15%, as stated on page 31. The most likely value, $m$, is informed from Figure 2, where 26% of 66 individuals reported using heroin on the day of discharge. Therefore, we estimated our model parameters using $\gamma=0$, $m=1$, and ${x}_{0.85}=30$ days to obtain $\mu_{F}=1.95$ and $\sigma_{F}=1.40$.
@@ A.6.7 Distribution (G) - Time in the inactive state given the previous state was active opioid use
Distribution (G) parameters were calibrated. The quantile ${x}_{q}$ was selected through model calibration by adding approximately $\pm$5from a starting value assessment that 90% of individuals used opioids within 20 years. The most likely value $m$ is selected through model calibration by adding approximately $\pm$5from a starting value 30. The final model parameters were estimated using $\gamma=0$, $m=1$, and ${x}_{0.7}=5.5*365.25$ days per year to obtain $\mu_{G}=6.29$ and $\sigma_{G}=2.51$.
@@ Appendix B Determining Number of Replications
$\hat{s}$ at $n=600$ 257.10 10.39 28.71 59.03 60.38
$h$ at $n=600$ 20.57 0.83 2.30 4.72 4.83
$\hat{n}$ 26 432 128 538 563
Additionally, we checked the residuals of the main outputs for normality by plotting and observing Q-Q plots and performing K-S tests of the main outputs. Table 8 reports the K-S test with Lilliefors correction and skewness. The K-S test fails to reject the null hypothesis that the number of arrests, number of hospital encounters, and number of treatments in 2032 follow a normal distribution. The K-S test rejects the null hypothesis for the number of individuals in the active use state and the number of opioid-related deaths at the end of 2032. Looking at the skewness in Table 8 we see that the active use state and opioid-related deaths skews slightly positively, meaning the mean estimates used in the main analysis are above the median of all $n$ runs. Figure 4 shows the Q-Q plots for the five main outputs of the model for years 2016, 2018, and 2032. Based on the relatively small skewness ($<0.75$) and Q-Q plots we conclude that all the main outputs are sufficiently normal to use confidence intervals as the main results, yet keep skewness in mind when interpreting the results. We also note that the normality assumption is valid for our cumulative results since the sum of two jointly normal random variables also follows a normal distribution.In fact, following the same process as above with cumulative outputs, showed the cumulative outputs required fewer than 400 replications to achieve a 95% CI. We have omitted the cumulative results due to their similarity to the yearly outputs.
@@
K-S Testα (p-value) 0.040 0.030 0.190 0.260 0.580
Skewness 0.240 0.060 0.020 0.140 0.70
@@ Appendix C Additional Simulated Results
Figure 7 illustrates the mean 95% CIs of arrests, hospitalizations, and treatment episodes per person over time. One is the lower bound of the re-arrest, re-hospitalization, and OUD treatment re-start rates. A value of one can be interpreted as everyone who used the given system only used it once in a given year. We expect the base model re-arrest and re-hospitalization rates to be a lower bound on the number of interactions for a given individual with these systems, since, as stated earlier, the model assumes“memoryless” cumulative state durations, whereas in reality some state transitions may be history-dependent (Nosyk et al.,, 2014). In Figure 7, we see statistically significant decreases in re-hospitalizations for policies with increases in OD, decreases in re-arrests for policies with AD or CM, and increases in OUD treatment for all policies in the first year or two after policy implementation. Figure 7 shows similar results to Table 5, but over time.
@@ Appendix D Sensitivity Analysis
This section explores and describes the sensitivity of the DES model. We conducted the sensitivity analysis on the scenario with arrest diversion at 60%, overdose diversion at 80%, and case management at 60%, referred to as scenario (60, 80, 60) for the remainder of this section. This scenario was selected for two main reasons. The first was due to the scenario’s high percentages of individuals diverted by the interventions. Two, the scenario’s scale of implementation is feasible for stakeholders to implement since we can expect some eligible individuals not to complete or participate in an available program (White et al.,, 2021). We evaluate changes to the following yearly outputs in 2016, 2018, and 2032: number of overdose deaths, individuals entering the CJS, hospital encounters, treatment, and active use. These outputs were selected since they were the main focus of the results. The outputs in the year 2016 test input sensitivity to the base model when no interventions have been implemented yet. The year 2018 indicates the sensitivity of the model with increased overdose diversion and arrest diversion implemented, and the year 2032 indicates the effects from the results still hold and the model sensitivity to when all three interventions are implemented.
@@
1 Warmup Initiation Age, $\mu$ 2.08 (1.98, 2.18)
2 Warmup Initiation Age, $\sigma$ 0.76 (0.72, 0.80)
3 Warmup Prevalence Age, $\mu$ 3.74 (3.55, 3.93)
4 Warmup Prevalence Age, $\sigma$ 0.49 (0.47, 0.51)
@@
Table 11 indicates three weak negative monotonic relationships between the inputs and outputs. The first is the relationship between the number of arrests and parameter 17, $\mu$ of the distribution for the time in the active state given the next event is an opioid-related arrest. The second is the relationship between the number of hospital encounters and parameter 15, $\mu$ of the distribution for the time in the active state given that the next event is an opioid-related hospital encounter. The third is the relationship between the number of OUD treatment starts and parameter 19, $\mu$ of the distribution for the time in the active state given the next event is OUD treatment. This relationship between OUD treatment and parameter 19 is less monotonic after implementing the OUD policies. As shown in Table 12, all three weak monotonicity relationships have effect sizes of -2.813, -1.385, and -2.409, respectively, the greatest of which is that of the relationship of parameter 17 and the number of opioid-related arrests. It is also worth noting the corresponding $\sigma$ parameters of these LN distributions had notable OLS effects of 1.154, 0.954, and -0.123, respectively, even though the relationships were very weakly monotone or determined non-monotone. This suggests that the corresponding model outputs are more sensitive to small changes to the $\mu$ parameters over the $\sigma$ parameters. Additionally, the adjustments are more predictable (i.e., linear) to the corresponding output with changes to the $\mu$ parameters. Therefore, parameters 15 through 20 should be carefully selected as they directly impact the magnitude of outputs over time. We do note that the effect size of parameter 15 on hospital encounters and parameter 18 on OUD treatment starts to decrease after the sustained implementation of the policies. This means that while they still impact model outputs, they seem to affect the outputs less after the policies are implemented.
@@
There are several other very weak monotone relationships in the model, such as between the number of opioid-related deaths and parameters 11, 12, 15, and 16. Parameter 11 is the $\mu$ of the distribution for the active state time given that the next event is an opioid-related death pre-2019. The corresponding $\sigma$ parameter, 12, is also very weakly monotone for this distribution. We note that the relationship between opioid-related deaths and parameter 13, the $\mu$ of the distribution for the active state time given that the next event is an opioid-related death post-2019, is non-monotone with a smaller effect size than that of parameter 11. Since the model is sensitive to parameters 11 and 12, care should be given to parameterizing opioid-related deaths. We also note that after the implementation of AD and OD, the absolute value of the OLS effects of parameters 11 and 12 reduces the magnitude from -1.527 to -0.23 and 1.365 to 0.184, respectively. This suggests that the AD and OD interventions make the model less sensitive to parameter 11. Parameters 15 and 16 are the $\mu$ and $\sigma$ of the distribution for the active state time given that the next event is a hospital encounter. Parameter 15 becomes more monotone, and its effect size increases after the AD and OD policies are implemented in 2018, but is decreased for 2032. This suggests that as parameter 15 increases, an OD policy would look more competitive for opioid-related deaths. The $\mu$ parameters 11, 13, and 15 have larger effect sizes than their corresponding distribution’s $\sigma$ parameters, meaning the model is more sensitive to the $\mu$ parameters of the distribution than their $\sigma$ parameters.
@@
Several additional very weakly monotone relationships exist for parameters 5, 19, 21, and 40. As the size of the total starting population (i.e., parameter 5) increases, the number of individuals in the active use state at the end of 2016 and 2018 increases. The effect reduces from 0.126 to 0.111 from 2016 to 2018 and is non-monotone with effect size 0.052 in 2032. This suggests that the starting population slightly affects the active use state, but its influence reduces the longer the simulation runs. Parameter 19, $\mu$ of the distribution for the time in the active state given the next event is OUD treatment, also has a very weak monotone relationship with the number in the active use state in 2032 with an effect size of 0.542. This means parameter 19 becomes more sensitive after implementing OUD treatment policies. Parameter 21, $\mu$ of the distribution for the time in the active state given the next event is to stop opioid use, is very weakly monotone, with the number of active use with an effect size of 0.675 in 2016, and OUD treatments with an effect size of 0.794 in 2018. Lastly, the $\mu$ of the LN distribution for the time in the inactive state following the active state (i.e., parameter 40) has a very weak monotone relationship with opioid-related deaths in 2032, opioid-related hospital encounters, the number of OUD treatments in 2018 and 2032 and the number of active use starts. This suggests that the model is most sensitive to parameter 40 as it impacts the most model outputs.
@@
The monotonic relationships between the input and output parameters described in this section are expected based on the model structure. Additionally, the effect size of parameters on outputs does not change as policies are implemented. This suggests that changes to these parameters would not affect the conclusions of this paper between the base model and treatment policy scenarios regarding opioid-related deaths, opioid-related arrests, hospital encounters, treatments, and active use.